# SignRoundV2: Closing the Performance Gap in Extremely Low-Bit Post-Training Quantization for LLMs

### v1

@@ Appendix A Detailed Data
Table 8: The detailed accuracies of 5 tasks for Llama2 models with INT2/4 mixed precision. "Avg. Bits" indicates average bit-width.
@@
Llama-2-7B 16Bits 16 - 42.83 75.97 57.27 77.91 69.30 64.66
@@
Table 9: The detailed accuracies of 5 tasks for Llama3 models with INT2/4 mixed precision. "Avg. Bits" indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. Piqa Wino. Avg.
Llama-3-8B 16Bits 16 - 50.17 80.09 60.14 79.54 73.24 68.64
AQLM 2 1x16 41.21 74.24 55.44 77.80 71.82 64.10
EfficientQAT 2 128 36.01 69.15 50.74 75.30 65.67 59.37
EfficientQAT 2 64 37.03 71.17 51.86 76.03 67.72 60.76
SignRoundV1 2 128 34.22 67.09 44.15 71.65 59.12 55.25
Ours 2 128 35.67 70.03 47.70 72.96 63.14 57.90
Ours 2 64 37.37 70.29 48.69 74.59 66.14 59.42
Ours* 2 128 38.99 72.18 49.19 73.94 65.57 59.97
Ours* 2 64 39.42 71.42 50.04 74.43 66.61 60.38
Ours 2.5 128 41.72 74.87 52.81 75.57 68.82 62.76
Ours 2.5 64 43.34 75.42 53.27 76.06 70.24 63.67
Ours* 2.5 128 41.38 74.87 53.31 75.46 67.96 62.60
Ours* 2.5 64 43.00 74.87 54.13 76.28 69.53 63.56
Llama-3-70B 16Bits 16 - 60.24 86.83 66.40 82.21 80.74 75.28
AQLM 2 1x16 50.34 78.83 63.47 79.65 78.22 70.10
EfficientQAT 2 128 48.81 79.25 60.75 79.60 69.46 67.57
EfficientQAT 2 64 49.06 77.40 61.60 77.37 74.03 67.89
SignRoundV1 2 128 47.87 79.71 43.96 78.62 73.64 64.76
Ours 2 128 49.91 80.81 59.75 79.49 75.14 69.02
Ours 2 64 52.22 81.73 60.93 80.09 75.69 70.13
Ours* 2 128 52.05 82.41 60.41 79.38 76.56 70.16
Ours* 2 64 54.95 82.53 61.05 79.82 76.32 70.93
Ours 2.5 128 56.83 83.67 63.06 80.47 77.82 72.37
Ours 2.5 64 55.29 83.08 62.60 81.23 77.66 71.97
Ours* 2.5 128 56.48 83.04 63.57 81.56 78.77 72.68
Ours* 2.5 64 56.40 82.95 63.07 80.63 78.53 72.32
Table 10: The detailed accuracies of 10 tasks for Llama-3.1-8B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
16 16Bits 51.54 81.78 84.04 59.13 73.14 67.96 33.00 80.03 37.21 73.80 64.16
4 RTN 44.37 75.55 80.86 55.19 61.96 57.06 31.80 76.28 29.25 70.80 58.31
DL 46.25 79.00 83.09 55.28 69.63 60.55 31.80 77.75 33.41 70.48 60.72
Ours 47.35 79.63 82.60 56.10 69.53 62.00 33.60 77.69 32.80 72.06 61.34
Ours 49.06 80.43 83.18 57.35 71.41 64.86 33.40 78.40 34.76 70.40 62.33
5 DL 50.43 79.17 84.10 58.22 68.62 64.77 34.60 78.62 34.52 72.61 62.57
Ours 49.83 80.93 84.19 57.92 72.19 65.43 35.80 79.22 35.37 71.03 63.19
6 DL 51.45 81.19 84.59 58.94 69.73 65.98 36.40 79.33 34.76 74.03 63.64
Ours 51.54 81.40 84.89 58.57 71.96 66.85 35.40 79.38 36.84 74.35 64.12
Table 11: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
16 16Bits 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.47 85.10 88.90 63.18 73.53 78.72 36.00 82.32 40.88 79.01 68.71
DL 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.90 85.82 88.53 63.63 75.45 79.97 37.00 81.88 40.76 80.27 69.32
Ours 59.64 86.24 88.69 64.51 76.89 81.20 37.20 82.86 40.39 82.00 69.96
5 DL 60.15 86.57 88.84 64.94 75.65 81.08 38.20 82.70 41.13 81.77 70.10
Ours 61.18 86.95 89.17 64.89 77.18 81.23 37.40 83.19 41.00 80.90 70.31
6 DL 61.09 86.53 88.93 65.45 76.89 81.99 38.00 83.08 41.13 82.48 70.56
Ours 62.12 86.15 88.72 65.65 76.69 82.00 37.60 83.30 40.88 81.85 70.50
Table 12: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
16 16Bits 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
DL 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 53.67 80.18 85.08 57.99 66.27 68.30 33.60 76.66 43.08 68.90 63.37
Ours 53.50 81.65 85.90 59.32 69.05 69.59 33.60 78.62 44.92 68.98 64.51
5 DL 51.11 80.26 86.85 59.54 68.27 69.42 35.20 76.77 44.31 68.67 64.04
Ours 54.61 82.62 85.90 59.78 69.67 70.43 33.60 78.07 45.90 69.77 65.04
6 DL 53.50 82.45 86.51 60.89 68.93 70.64 36.80 77.91 46.39 70.48 65.45
Ours 54.18 82.24 86.30 60.68 69.73 70.52 34.80 78.73 45.78 70.01 65.30
Table 13: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
16 16Bits 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
DL 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.84 80.98 85.84 53.09 61.91 68.88 32.00 75.30 36.35 70.72 61.89
Ours 53.75 81.61 86.21 54.26 64.16 70.25 31.40 75.57 37.09 68.51 62.28
5 DL 51.11 80.30 86.09 54.23 62.04 70.26 31.40 75.19 35.62 68.27 61.45
Ours 54.10 81.82 86.42 54.62 64.20 71.53 30.80 76.01 37.58 68.27 62.54
6 DL 53.84 81.57 86.94 55.36 62.10 71.46 31.20 75.95 35.50 66.06 62.00
Ours 52.65 81.65 86.97 55.57 64.06 71.54 30.20 76.17 35.74 67.32 62.19
Table 14: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/8 mixed precision. "Avg. Bits" indicates average per-weight bit-width. "DL" means DeltaLoss-only without tuning.
16 16Bits 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 55.38 81.82 85.54 61.54 65.67 76.39 34.40 78.94 38.80 72.22 65.07
DL 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 59.47 83.16 87.19 61.70 68.78 78.39 35.40 79.71 41.74 73.01 66.86
Ours 57.59 84.09 88.23 62.51 69.42 79.51 33.00 79.49 41.00 74.03 66.89
5 DL 56.83 83.38 84.50 63.18 66.25 79.58 34.00 79.92 40.39 71.98 66.00
Ours 57.85 84.47 87.92 63.33 67.22 80.10 35.60 80.25 40.51 74.43 67.17
6 DL 57.51 83.04 86.12 63.28 67.69 80.02 36.40 80.36 40.15 71.51 66.61
Ours 58.19 83.88 86.85 63.44 68.12 80.40 36.40 80.47 40.88 73.48 67.21

### v2

@@ C.2 Mixed MXFP4/8 Results
Table 20: The detailed accuracies of 10 tasks for Llama-3.1-70B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
16 16-bit 62.46 86.87 87.80 65.15 75.51 82.34 37.20 83.30 40.64 78.69 70.00
4 RTN 59.47 85.10 88.90 63.18 73.53 78.72 36.00 82.32 40.88 79.01 68.71
SRV1 58.28 85.06 88.53 63.35 75.24 79.66 37.60 82.32 40.15 79.95 69.01
Ours 59.90 85.82 88.53 63.63 75.45 79.97 37.00 81.88 40.76 80.27 69.32
Ours 59.64 86.24 88.69 64.51 76.89 81.20 37.20 82.86 40.39 82.00 69.96
5 DL 60.15 86.57 88.84 64.94 75.65 81.08 38.20 82.70 41.13 81.77 70.10
Ours 61.18 86.95 89.17 64.89 77.18 81.23 37.40 83.19 41.00 80.90 70.31
6 DL 61.09 86.53 88.93 65.45 76.89 81.99 38.00 83.08 41.13 82.48 70.56
Ours 62.12 86.15 88.72 65.65 76.69 82.00 37.60 83.30 40.88 81.85 70.50
Table 21: The detailed accuracies of 10 tasks for Qwen2.5-7B-Instruct model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
16 16-bit 52.56 81.61 86.36 62.01 69.44 71.76 34.60 79.65 47.98 70.72 65.67
4 RTN 48.81 75.59 82.48 56.84 57.68 65.15 34.80 75.79 41.62 67.48 60.62
SRV1 53.24 80.89 85.47 57.88 64.41 68.05 35.00 77.48 46.76 71.43 64.06
Ours 53.67 80.18 85.08 57.99 66.27 68.30 33.60 76.66 43.08 68.90 63.37
Ours 53.50 81.65 85.90 59.32 69.05 69.59 33.60 78.62 44.92 68.98 64.51
5 DL 51.11 80.26 86.85 59.54 68.27 69.42 35.20 76.77 44.31 68.67 64.04
Ours 54.61 82.62 85.90 59.78 69.67 70.43 33.60 78.07 45.90 69.77 65.04
6 DL 53.50 82.45 86.51 60.89 68.93 70.64 36.80 77.91 46.39 70.48 65.45
Ours 54.18 82.24 86.30 60.68 69.73 70.52 34.80 78.73 45.78 70.01 65.30
Table 22: The detailed accuracies of 10 tasks for Qwen3-8B model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
16 16-bit 55.46 83.38 86.73 57.04 64.00 72.96 31.20 76.61 36.72 68.27 63.24
4 RTN 50.17 76.14 84.98 52.00 58.08 65.90 28.60 73.01 33.66 62.83 58.54
SRV1 51.28 81.02 85.20 52.86 58.26 68.51 31.40 75.03 32.44 66.54 60.25
Ours 53.84 80.98 85.84 53.09 61.91 68.88 32.00 75.30 36.35 70.72 61.89
Ours 53.75 81.61 86.21 54.26 64.16 70.25 31.40 75.57 37.09 68.51 62.28
5 DL 51.11 80.30 86.09 54.23 62.04 70.26 31.40 75.19 35.62 68.27 61.45
Ours 54.10 81.82 86.42 54.62 64.20 71.53 30.80 76.01 37.58 68.27 62.54
6 DL 53.84 81.57 86.94 55.36 62.10 71.46 31.20 75.95 35.50 66.06 62.00
Ours 52.65 81.65 86.97 55.57 64.06 71.54 30.20 76.17 35.74 67.32 62.19
Table 23: The detailed accuracies of 10 tasks for Qwen3-32B model with MXFP4/8 mixed precision. “Avg. Bits” indicates average per-weight bit-width. “DL” denotes DeltaLoss-only without tuning.
16 16-bit 57.94 84.47 86.39 63.91 67.13 80.74 36.00 80.96 39.05 73.40 67.00
4 RTN 55.38 81.82 85.54 61.54 65.67 76.39 34.40 78.94 38.80 72.22 65.07
SRV1 57.51 84.55 87.13 61.61 68.83 78.32 36.20 79.71 41.37 73.95 66.92
Ours 59.47 83.16 87.19 61.70 68.78 78.39 35.40 79.71 41.74 73.01 66.86
Ours 57.59 84.09 88.23 62.51 69.42 79.51 33.00 79.49 41.00 74.03 66.89
5 DL 56.83 83.38 84.50 63.18 66.25 79.58 34.00 79.92 40.39 71.98 66.00
Ours 57.85 84.47 87.92 63.33 67.22 80.10 35.60 80.25 40.51 74.43 67.17
6 DL 57.51 83.04 86.12 63.28 67.69 80.02 36.40 80.36 40.15 71.51 66.61
Ours 58.19 83.88 86.85 63.44 68.12 80.40 36.40 80.47 40.88 73.48 67.21
Appendix D NVFP4 Results
This section provides detailed NVFP4 accuracy results reported in Table 24. Our method achieves the best average accuracy on four out of five models and remains competitive on Qwen3-32B.
Table 24: Comparison of SignRoundV1 (SRV1), RTN, and our method under NVFP4 quantization. We report accuracy (%) across 10 tasks (listed in Section 4.1) and the average (AVG). The “I” in the model name is short for "Instruct".
Model Method ARC-c. ARC-e BoolQ Hella. Lamb. MMLU Open. PIQA Truth. Wino. AVG
Llama3.1-8B-I 16-bit 53.75 82.15 85.35 59.79 72.06 68.28 35.60 80.30 38.07 73.32 64.87
RTN 50.60 80.01 83.43 58.30 71.34 64.13 35.00 78.89 34.27 71.90 62.79
SRV1 50.26 80.39 84.28 58.23 71.07 65.39 34.00 79.22 35.86 72.77 63.15
Ours 50.68 81.65 85.44 57.93 71.49 64.91 35.00 79.49 37.58 74.51 63.87
Llama3.1-70B-I 16-bit 61.09 86.20 89.11 66.28 76.46 82.52 38.80 83.73 41.98 83.43 70.96
RTN 60.58 85.94 89.45 65.03 75.47 80.54 37.20 83.24 42.11 81.14 70.07
SRV1 59.73 85.98 88.84 64.77 75.59 81.31 39.00 83.13 40.88 82.32 70.16
Ours 60.75 86.87 88.87 65.10 75.82 81.25 38.20 82.54 41.00 81.29 70.17
Qwen2.5-7B-I 16-bit 53.07 81.69 86.33 62.06 69.38 71.71 34.60 79.71 47.98 71.03 65.76
RTN 51.96 80.18 86.24 60.65 67.15 70.47 33.20 78.51 43.82 67.96 64.01
SRV1 52.82 81.14 86.09 60.90 69.07 70.61 34.80 78.67 47.61 70.40 65.21
Ours 54.18 81.69 86.79 60.23 69.65 70.67 34.00 78.51 47.12 70.96 65.38
Qwen3-8B 16-bit 55.20 83.54 86.61 57.16 64.10 72.90 31.60 76.66 36.47 67.96 63.22
RTN 53.41 82.37 85.87 55.29 62.45 70.82 30.80 76.44 33.66 67.56 61.87
SRV1 54.86 82.91 86.88 55.31 62.95 71.97 31.20 76.28 36.35 67.48 62.62
Ours 55.29 83.04 86.85 55.41 63.71 71.36 29.20 75.79 37.94 69.22 62.78
Qwen3-32B 16-bit 58.02 84.39 86.42 63.89 67.13 80.74 36.20 81.23 38.56 73.09 66.97
RTN 57.51 84.01 87.03 63.39 65.83 79.80 35.40 79.43 39.53 71.51 66.34
SRV1 59.56 84.22 88.01 62.99 68.29 79.50 37.60 79.54 41.00 72.22 67.29
Ours 57.51 84.68 86.51 63.09 68.89 79.80 36.00 79.60 39.05 72.38 66.75
Appendix E INT2/4 Mixed-bit Results
This appendix provides the detailed task-wise results for the INT2/4 mixed-bit weight-only quantization experiments discussed in Section 4.2. We report accuracies on ARC-Challenge, ARC-Easy, HellaSwag, PIQA, and WinoGrande for both uniform and mixed-precision settings across different model scales. These results complement the averaged comparisons in Table 1 and provide a more fine-grained view of the behavior of each method under extreme low-bit quantization. Specifically, Table 25 reports the detailed accuracies for Llama-3 models, and Table 26 reports the results for Llama-2 models.
Table 25: The detailed accuracies of 5 tasks for Llama3 models with INT2/4 mixed precision. “Avg. Bits” indicates average bit-width.
@@ Appendix E INT2/4 Mixed-bit Results
Llama3-8B 16-bit 16 - 50.17 80.09 60.14 79.54 73.24 68.64
AQLM 2 1x16 41.21 74.24 55.44 77.80 71.82 64.10
EfficientQAT 2 128 36.01 69.15 50.74 75.30 65.67 59.37
EfficientQAT 2 64 37.03 71.17 51.86 76.03 67.72 60.76
SignRoundV1 2 128 34.22 67.09 44.15 71.65 59.12 55.25
Ours 2 128 35.67 70.03 47.70 72.96 63.14 57.90
Ours 2 64 37.37 70.29 48.69 74.59 66.14 59.42
Ours* 2 128 38.99 72.18 49.19 73.94 65.57 59.97
Ours* 2 64 39.42 71.42 50.04 74.43 66.61 60.38
Ours 2.5 128 41.72 74.87 52.81 75.57 68.82 62.76
Ours 2.5 64 43.34 75.42 53.27 76.06 70.24 63.67
Ours* 2.5 128 41.38 74.87 53.31 75.46 67.96 62.60
Ours* 2.5 64 43.00 74.87 54.13 76.28 69.53 63.56
Llama3-70B 16-bit 16 - 60.24 86.83 66.40 82.21 80.74 75.28
AQLM 2 1x16 50.34 78.83 63.47 79.65 78.22 70.10
EfficientQAT 2 128 48.81 79.25 60.75 79.60 69.46 67.57
EfficientQAT 2 64 49.06 77.40 61.60 77.37 74.03 67.89
SignRoundV1 2 128 47.87 79.71 43.96 78.62 73.64 64.76
Ours 2 128 49.91 80.81 59.75 79.49 75.14 69.02
Ours 2 64 52.22 81.73 60.93 80.09 75.69 70.13
Ours* 2 128 52.05 82.41 60.41 79.38 76.56 70.16
Ours* 2 64 54.95 82.53 61.05 79.82 76.32 70.93
Ours 2.5 128 56.83 83.67 63.06 80.47 77.82 72.37
Ours 2.5 64 55.29 83.08 62.60 81.23 77.66 71.97
Ours* 2.5 128 56.48 83.04 63.57 81.56 78.77 72.68
Ours* 2.5 64 56.40 82.95 63.07 80.63 78.53 72.32
Table 26: The detailed accuracies of 5 tasks for Llama2 models with INT2/4 mixed precision. “Avg. Bits” indicates average bit-width.
Model Method Avg. Bits Group Size ARC-c ARC-e Hella. PIQA Wino. Avg.
Llama2-7B 16-bit 16 - 42.83 75.97 57.27 77.91 69.30 64.66